### v1

# Better Safe Than Sorry? Overreaction Problem of Vision Language Models in Visual Emergency Recognition

Abstract
Vision-Language Models (VLMs) have demonstrated impressive capabilities in understanding visual content, but their reliability in safety-critical contexts remains under-explored. We introduce VERI (Visual Emergency Recognition Dataset), a carefully designed diagnostic benchmark of 200 images (100 contrastive pairs). Each emergency scene is matched with a visually similar but safe counterpart through multi-stage human verification and iterative refinement. Using a two-stage protocol—risk identification and emergency response—we evaluate 14 VLMs (2B-124B parameters) across medical emergencies, accidents, and natural disasters. Our analysis reveals a systematic overreaction problem: models excel at identifying real emergencies (70-100% success rate) but suffer from an alarming rate of false alarms, misidentifying 31-96% of safe situations as dangerous, with 10 scenarios failed by all models regardless of scale. This "better-safe-than-sorry" bias manifests primarily through contextual overinterpretation (88-93% of errors), challenging VLMs’ reliability for safety applications. These findings highlight persistent limitations that are not resolved by increasing model scale, motivating targeted approaches for improving contextual safety assessment in visually misleading scenarios.
[…]
tcb@breakable
Better Safe Than Sorry? Overreaction Problem of Vision Language Models in Visual Emergency Recognition
Dasol Choi1,2  Seunghyun Lee1  Youngsook Song3thanks: Corresponding Author Yonsei University1  MODULABS2  Lablup Inc.3 dasolchoi@yonsei.ac.kr  lutris@yonsei.ac.kr  yssong@lablup.com
Content Warning: This paper contains images and descriptions of emergency situations.
Figure 1: The overreaction problem in VLMs: correctly identifying actual emergencies (left) while misclassifying visually similar safe scenarios as dangerous (right), similar to human misperceptions of TV fire videos as real fires in 2023.
Figure 2: Examples from the VERI dataset showing contrastive pairs across three categories. Top row (red background): genuine emergency situations. Bottom row (green background): visually similar but safe scenarios. Each pair maintains visual similarity while representing different semantic meanings—one requiring intervention and the other representing safe activities. These pairs enable evaluation of VLMs’ ability to make safety distinctions despite visual similarities.
1 Introduction
Vision-Language Models (VLMs) have made remarkable progress in understanding visual content, advancing from simple object recognition to sophisticated scene understanding and contextual reasoning (OpenAI, 2023; Google, 2023; Anthropic, 2024). These models now power a wide range of applications from content moderation to assistive technologies (Zhang et al., 2024). However, their reliability in safety-critical scenarios remains underexplored. This study investigates a critical question with real-world implications: To what extent can current VLMs distinguish between genuine emergencies and visually similar but safe scenarios?
False visual perceptions of emergencies can lead to costly resource mobilization. In October 2023, firefighters were dispatched to two separate incidents in New York and Seoul. In both cases, they discovered that high-definition fireplace videos playing on television screens had been mistaken for actual fires (Lloyd, 2023; Yoon, 2023). Such false alarms carry substantial costs. Each incident costs approximately \$1,000-2,500, with annual costs exceeding \$1 billion in the United States alone (Square, 2024). As VLMs power applications from smart home monitoring to CCTV surveillance, understanding their tendency to "overreact" becomes crucial (Yun and Lim, 2025).
Current research on safety-critical VLMs has primarily focused on specialized domains. These include medical diagnosis systems (Ben Abacha et al., 2019), autonomous driving perception (Caesar et al., 2020), and industrial robotics safety mechanisms (Dhillon et al., 2002). Our work addresses a distinct gap: everyday emergency recognition across domains. This task requires different contextual reasoning capabilities—distinguishing between visually similar scenarios based on subtle contextual cues rather than domain-specific features. Existing benchmarks typically evaluate models on isolated images, rarely using contrastive pairs that test the ability to distinguish between visually similar but semantically distinct situations.
To address these gaps, we introduce VERI (Visual Emergency Recognition Dataset) [^1] The VERI dataset is publicly available at
https://huggingface.co/datasets/Dasool/VERI-Emergency
, a diagnostic benchmark of 200 images comprising 100 contrastive pairs. Inspired by compact yet influential diagnostic benchmarks such as Winogender (720 examples) (Rudinger, 2018), WSC (60 examples) (Levesque et al., 2012), and PAIRS (200 images) (Fraser and Kiritchenko, 2024), VERI prioritizes diagnostic value over quantity. We employ controlled image pairs to evaluate VLMs’ ability to distinguish between emergency and non-emergency situations. Each pair underwent over 5 rounds of iterative refinement and verification to balance visual similarity with semantic distinction.
Our work makes the following contributions:
  • We introduce VERI, a benchmark with a two-stage protocol assessing risk identification and action suggestion in emergency scenarios.
  • Through evaluation of 14 VLMs (2B-124B parameters), we reveal a systematic "overreaction problem" persisting across architectures and scales, with 10 safe scenarios misclassified by all models.
  • We identify two error patterns (contextual and visual misinterpretations), with contextual errors dominating (88-93%) and varying by emergency category.
  • We show the overreaction problem persists despite model scaling, requiring targeted architectural improvements for contextual emergency assessment.
Category Accidents & Unsafe Behaviors Personal Medical Emergencies Natural Disasters
Scope Immediate physical dangers from environment or human action Urgent health risks to individuals Large-scale threats affecting multiple people
Example scenarios Traffic accidents, falls from heights, drowning risks, physical altercations, unsafe tool use Cardiac arrest, choking, unconsciousness, severe injuries, allergic reactions, seizures Fires, floods, earthquakes, building collapses, hurricanes, avalanches
Table 1: Taxonomy and examples of emergency situations in the VERI dataset.
2 VERI: Visual Emergency Recognition Dataset
Effective emergency detection requires distinguishing between genuine threats and visually similar but safe situations. To evaluate this capability in Vision Language Models, we introduce VERI (Visual Emergency Recognition Dataset), a benchmark designed to challenge models’ contextual understanding in safety-critical scenarios.
2.1 Dataset Design and Taxonomy
The core design principle of VERI is the contrastive pair approach—each entry consists of two images with high visual similarity but fundamentally different semantic implications: one depicting a genuine emergency requiring intervention and the other showing a visually similar but safe scenario. As illustrated in Figure 2, these pairs include contrasts such as actual medical emergencies versus training simulations, genuine accidents versus staged scenarios, and real disasters versus controlled environments. This contrastive structure enables direct assessment of a model’s ability to distinguish between superficially similar scenes with different safety implications.
We organized VERI into three major categories of emergency situations (Accidents & Unsafe Behaviors, Personal Medical Emergencies, and Natural Disasters), as shown in Table 1. Each category presents unique visual challenges, from subtle physiological distress cues to complex environmental interpretation, as demonstrated in Figure 2.
2.2 Image Creation Process
Creating visually similar but semantically distinct image pairs required careful design to balance realism with controlled visual properties. We employed a multi-stage approach:
Stage 1: Scenario Definition. For each category, we defined specific scenario pairs (e.g., "person requiring CPR" vs. "CPR training on mannequin"), identifying visual elements to preserve or alter for semantic distinction.
Stage 2: Image Generation. Two researchers created prompts for image generation using GPT-4o. We preserved key visual elements while altering critical semantic cues (e.g., real person vs. mannequin, distress vs. educational context).
Stage 3: Refinement and Validation. We iteratively refined images using inpainting until they matched our requirements (averaging 5+ cycles per pair). Three evaluators (two researchers and an independent annotator) validated each image against our criteria. The process continued until unanimous agreement was reached on all criteria (Cohen’s $\kappa=1.00$, $n=200$ pairs, $600$ decisions), eliminating annotation noise. These criteria ensured emergency scenes clearly required intervention, safe scenes were unambiguously non-threatening, and pairs maintained appropriate visual similarity.
Stage 4: Caption Generation. We created descriptive captions for each image pair using GPT-4o, which were then refined to ensure accuracy.
Our process prioritized diagnostic value over quantity, focusing on expert-guided refinement to effectively challenge VLMs’ contextual reasoning.
2.3 Query Development
Our evaluation protocol requires two types of queries to assess different aspects of model performance:
Q1: Risk Identification. The first query—"Does the situation in this image appear dangerous or safe?"—is a straightforward binary classification task designed to evaluate models’ ability to distinguish between dangerous and safe situations.
Q2: Emergency Response. For emergency images, we used the query—"Based on this situation, what emergency response should be taken?"—to assess models’ ability to recommend appropriate actions. We created a standardized response bank through a three-step process. First, we used image captions as input to Perplexity to generate initial responses based on authoritative sources. Next, three evaluators verified these responses against original sources. Finally, we compiled critical actions for each emergency type as evaluation criteria.
Statistic Count
Total image pairs 100
Total images 200
Accidents & Unsafe Behaviors pairs 35
Personal Medical Emergencies pairs 33
Natural Disasters pairs 32
Risk identification QA (Q1) 200
Emergency Response QA (Q2) 100
Detailed image captions 200
Table 2: VERI dataset statistics
Model Q1: Risk Identification Q2: Emergency Response
Precision Recall F1 Score # Images
Qwen2.5-VL Family
Qwen2.5-VL (3B) $0.510$ $1.000$ $0.676$ $0.460$ 98
Qwen2.5-VL (7B) $0.554$ $0.880$ $0.680$ $0.618$ 88
Qwen2.5-VL (32B) $0.589$ $0.890$ $0.709$ $0.6972$ 88
Qwen2.5-VL (72B) $0.652$ $0.900$ $0.756$ $0.7000$ 89
LLaVA-Next Family
LLaVA-Next (7B) $0.577$ $0.970$ $0.724$ $0.466$ 95
LLaVA-Next (13B) $0.575$ $1.000$ $0.730$ $0.502$ 98
InternVL3 Family
InternVL3 (2B) $0.633$ $0.950$ $0.760$ $0.497$ 93
InternVL3 (8B) $0.721$ $0.800$ $0.758$ $0.610$ 80
InternVL3 (14B) $0.658$ $0.960$ $0.781$ $0.638$ 94
Mistral Family
Mistral-Small (24B) $0.572$ $0.950$ $0.714$ $0.625$ 93
Pixtral (12B) $0.654$ $0.890$ $0.754$ $0.594$ 89
Pixtral-Large (124B) $0.632$ $0.980$ $0.769$ $0.677$ 96
Other Models
Idefics2 (8B) $0.528$ $0.950$ $0.679$ $0.463$ 93
Phi-3.5-vision (4B) $0.620$ $0.700$ $0.657$ $0.471$ 70
Table 3: Performance evaluation across risk identification (Q1) and emergency response (Q2) tasks. Q1 metrics show models’ ability to distinguish between dangerous and safe situations. Q2 scores reflect the quality of suggested actions for correctly identified emergencies, with # Images indicating the number of emergency images for which the model provided recommendations.
2.4 Dataset Statistics
The final VERI dataset consists of 100 image pairs (200 total images) distributed across three categories: Accidents & Unsafe Behaviors (35 pairs), Personal Medical Emergencies (33 pairs), and Natural Disasters (32 pairs). For evaluation, we created 200 binary classification questions (Q1) covering all images and 100 open-ended response questions (Q2) for emergency images only. Each image is accompanied by a detailed caption describing the scene context, key elements, and situational details, providing additional textual information that can be used for multimodal training or analysis. Figure 2 illustrates representative examples from each category, demonstrating both the visual similarity within pairs and the semantic distinction between emergency and safe scenarios.
3 Experimental Settings
3.1 Models
We evaluated 14 open-source VLMs (2B-124B parameters) across different architectures: Qwen2.5-VL (Bai et al., 2025) (3B, 7B, 32B, 72B), transformer-based models optimized for visual recognition; LLaVA-Next (Liu et al., 2024) (7B, 13B), integrating CLIP encoders with large language models; InternVL3 (Zhu et al., 2025) (2B, 8B, 14B), using a "ViT-MLP-LLM" architecture; the Mistral family , including Mistral-Small (24B) (AI, 2025) and Pixtral variants (12B, 124B) (Agrawal et al., 2024); and other architectures represented by Idefics2 (8B) (Laurençon et al., 2024) and Phi-3.5-vision (4B) (Abdin et al., 2024). All models were evaluated using publicly available checkpoints without task-specific fine-tuning to assess zero-shot capabilities.
(a) Precision-Recall tradeoff
(b) False Positive vs False Negative rates
Figure 3: Performance analysis of VLMs on emergency detection tasks. (a) Shows the pattern of high recall but lower precision across models, with point size indicating model parameter count. (b) Reveals a consistent “better-safe-than-sorry” bias where safe images are misclassified more frequently than emergencies are missed.
3.2 Evaluation Protocol
We evaluate model responses using a two-stage protocol:
Q1 (Binary Classification). The model is asked to answer whether the situation in the image is dangerous or safe. It selects one of two choices (A. Dangerous / B. Safe) and provides a brief reasoning. We compute precision, recall, and F1 score based on human-annotated binary ground truth labels.
Q2 (Open-ended Response). For images classified as dangerous by the model and correctly aligned with the ground truth, we further ask what an appropriate emergency response should be. We evaluated these responses using GPT-4o as a judge, providing it with the image caption and our curated gold-standard answers as references. The judge was instructed to score each response on a scale from 0 to 1 based on its alignment with the reference materials and appropriateness for the emergency situation.
4 Results and Analysis
4.1 The Overreaction Phenomenon
To evaluate VLMs’ emergency recognition capabilities, we assessed their performance on both risk identification (Q1) and emergency response (Q2) tasks (Table 3). Our evaluation reveals a consistent pattern: models achieve high recall (0.70-1.00) in identifying dangerous situations, but precision is notably lower (0.51-0.72), indicating a systematic "better-safe-than-sorry" bias—an "overreaction problem." This pattern persists across model families and parameter scales (2B-124B). Model size does not consistently correlate with improved precision; InternVL3 (8B) achieves the highest precision (0.72), outperforming larger variants. False positive rates (safe images misclassified as dangerous) ranged from 31% to 96%—substantially higher than false negative rates (missed emergencies), which ranged from 0% to 30%. Figure 3(a) illustrates this bias through the precision-recall tradeoff, with models clustering in the high-recall but lower-precision region.
Particularly concerning is that 10% of safe scenarios were misclassified by all 14 evaluated models, typically containing visual elements strongly associated with danger despite clear contextual safety cues. As shown in Figure 3(b), false positives consistently outnumber false negatives across all models. Increasing model size did not reliably mitigate this problem—within the Qwen2.5-VL family, precision improves modestly from 3B (0.51) to 72B (0.65), but recall fluctuates non-linearly. This persistent pattern across architectures and scales suggests the overreaction problem may be embedded in foundational visual understanding rather than reasoning capacity.
Figure 4: Error patterns in danger assessment: Visual Misinterpretation (blue) vs. Contextual Overinterpretation (red). Examples show how VLMs incorrectly classify safe situations as dangerous through either misperceiving visual elements or exaggerating potential risks in properly perceived scenarios.
4.2 Emergency Response Evaluation
Beyond identifying emergencies, we evaluated how effectively models suggest appropriate actions for emergency situations. Models demonstrated moderate effectiveness (scores 0.46-0.70) in generating appropriate emergency responses for correctly identified dangerous scenarios.
Unlike risk identification, emergency response quality shows a clearer correlation with model size. We observe consistent scaling benefits: Qwen2.5-VL (0.46→0.70), InternVL3 (0.50→0.64), and Mistral (0.59→0.68). This linear relationship presents an interesting contrast to the overreaction problem: while contextual reasoning for risk assessment doesn’t consistently improve with scaling, procedural knowledge for emergency responses does. This suggests that protocols may be more amenable to parameter scaling than the nuanced contextual judgments required for accurate risk identification.
We found a disconnect between risk identification and response capabilities. InternVL3 (8B) achieved the highest precision in risk identification but only moderate response quality, while models with lower precision often produced better emergency responses. This suggests VLMs develop these capabilities through different mechanisms—procedural knowledge scaling more predictably than contextual reasoning. Even when models correctly identify emergencies, their ability to recommend appropriate interventions remains limited. Examples can be found in Appendix A.3.
4.3 Category-Specific Analysis
Our analysis reveals measurable performance variations across emergency categories (Table 4). In risk identification (Q1), models generally performed best on Natural Disasters (ND, avg F1=0.75) compared to Accidents & Behaviors (AB, 0.73) and Personal Medical Emergencies (PME, 0.68). The PME category shows the highest variance, from exceptional (0.79 for Pixtral-Large) to extremely poor (0.41 for Phi-3.5-vision).
Model scaling effects vary by category. For PME, larger models generally show improved performance (e.g., Qwen2.5-VL improves from 0.68 to 0.73, Pixtral from 0.67 to 0.79), suggesting increased parameters help with subtle physiological cues. For AB and ND categories, we observe a non-linear relationship with size, where mid-sized models sometimes outperform larger counterparts (e.g., InternVL3-8B achieves 0.83 on AB, outperforming the 14B variant). The overreaction problem is most pronounced in the PME category (38% false positive rate) where models frequently misclassified training scenarios as genuine emergencies. Natural Disasters show the lowest false positive rate (27%), likely because environmental containment features provide more distinctive safety indicators.
These findings suggest VLMs struggle most with scenarios where emergency status depends on fine-grained contextual details rather than obvious visual elements. For emergency response quality, Q1 showed a clearer category preference (ND > AB > PME with differences up to 0.07), while Q2 exhibited smaller average differences between categories (approximately 0.03). This suggests that once a model correctly identifies an emergency, its ability to generate appropriate responses is more consistent across different emergency types. For detailed analysis of model-specific performance across categories, see Appendix A.2.
Model PME AB ND
Qwen2.5-VL Family
Qwen2.5-VL (3B) 0.681 0.673 0.674
Qwen2.5-VL (7B) 0.583 0.737 0.696
Qwen2.5-VL (32B) 0.712 0.697 0.719
Qwen2.5-VL (72B) 0.727 0.764 0.771
LLaVA-Next Family
LLaVA-Next (7B) 0.729 0.695 0.750
LLaVA-Next (13B) 0.727 0.729 0.733
InternVL3 Family
InternVL3 (2B) 0.753 0.747 0.781
InternVL3 (8B) 0.593 0.825 0.805
InternVL3 (14B) 0.767 0.761 0.815
Mistral Family
Mistral-Small (24B) 0.700 0.707 0.736
Pixtral (12B) 0.667 0.767 0.815
Pixtral-Large (124B) 0.790 0.745 0.777
Other Models
Idefics2 (8B) 0.683 0.693 0.660
Phi-3.5-vision (4B) 0.408 0.707 0.756
Table 4: F1 scores across different emergency categories (PME: Personal Medical Emergencies, AB: Accidents & Behaviors, ND: Natural Disasters)
4.4 Error Patterns and Analysis
Our analysis reveals two primary patterns in the overreaction problem (false positives), where models misclassify safe situations as dangerous, as illustrated in Figure 4:
Visual Misinterpretation: Models incorrectly perceive visual elements, failing to distinguish between safe and dangerous scenarios (e.g., confusing mannequins with real people, theatrical makeup with real injuries).
Contextual Overinterpretation: Models correctly identify visual elements but fail to properly interpret their safety implications within the broader context. This manifests as an exaggeration of potential risks by misweighing contextual factors. For example, models claim "the child’s shirt could lead to choking hazards" or "the car’s rearview mirror could blind the cyclist.
Our analysis across 14 models revealed that Contextual Overinterpretation was dominant, accounting for 88-93% of all misclassifications regardless of model architecture or scale. Notably, in the Natural Disasters category, 100% of the errors were Contextual Overinterpretation, suggesting that models can correctly identify elements like fire or water but consistently fail to assess their contextual safety.
These findings suggest that current VLMs can detect potentially hazardous elements, but lack the nuanced reasoning required to assess whether these elements pose actual dangers in specific contexts—a critical limitation for their reliability in safety applications. For detailed error pattern analysis across model sizes and categories, see Appendix C, and additional examples of Contextual Overinterpretation can be found in Appendix B.
4.5 Model Size and Category Effects
Our analysis reveals a non-linear relationship between model size and performance (Kaplan et al., 2020). The InternVL3 family exhibits a "Goldilocks effect," (Miller and Alexander, 2025) where the mid-sized 8B model excels in Accidents & Behaviors (0.83) and Natural Disasters (0.81), yet underperforms in Medical Emergencies (0.59) compared to both smaller 2B (0.75) and larger 14B (0.77) variants (see Table 4).
This suggests categories with obvious visual danger cues benefit from mid-sized models balancing perception and reasoning. Conversely, medical emergencies—requiring finer distinctions—perform better with either very small models (efficient pattern-matching) or large models (enhanced reasoning). The precision-recall tradeoff varies inconsistently with scaling. In Qwen2.5-VL, precision improves modestly with size (0.51→0.56→0.57), but recall fluctuates (1.00→0.88→1.00), challenging assumptions about model scaling in safety tasks.
Contextual Overinterpretation dominated across all parameter scales (89-91%). Even the best-performing model (InternVL3-8B) showed similar patterns (90.3% Contextual Overinterpretation), indicating a fundamental limitation in contextual reasoning that persists regardless of size. These findings suggest emergency recognition requires specialized architectures or fine-tuning approaches tailored to each category’s unique challenges.
Figure 5: Media-based danger misclassification. Left: Drive-in theater thunderstorm scene. Right: Flood poster viewed by pedestrian. Both cases show models failing to recognize representational contexts.
Figure 6: Visual similarity misclassification. Left: Ketchup mistaken for blood. Right: Training mannequin or staged combat confused with real danger.
4.6 Universal Misclassifications
Most striking is that 10 of our 100 safe images (10%) were misclassified by all 14 evaluated models, revealing common triggers for overreaction across architectures and parameter scales. These universally misclassified images represent the most extreme cases of the overreaction problem, with Visual Misinterpretation being the dominant factor.
The Visual Misinterpretation errors manifested in two key ways. First, models consistently failed to recognize representational contexts, as shown in Figure 5, where dangerous elements were portrayed in media rather than occurring in reality. Models misclassified scenes showing thunderstorms on drive-in theater screens or flood imagery on posters because they failed to perceive the critical visual cues that indicated these were representations. Second, as illustrated in Figure 6, models confused visually similar but contextually distinct scenarios, such as mistaking ketchup for blood or training mannequins for real people in danger.
5 Discussion
5.1 Balancing Safety and Accuracy
Our findings highlight a fundamental tension in visual emergency recognition: the trade-off between sensitivity to potential dangers and precision in distinguishing genuine emergencies from safe scenarios. While high sensitivity is desirable, the substantial false positive rates we observed (28-48%) could undermine practical utility in real-world applications. False alarms in emergency services lead to resource misallocation and "alarm fatigue," where users begin to distrust system warnings. Conversely, increasing precision must consider the severe consequences of missed emergencies. This optimization challenge varies across domains - medical triage might justifiably prioritize recall, while home monitoring systems require higher precision for user trust. The higher false positive rates in medical emergencies (38%) compared to natural disasters (27%) further suggest different domains require tailored approaches to balancing sensitivity and specificity.
5.2 Implications for Model Development
Our findings suggest several approaches for improving VLMs’ emergency recognition capabilities. Enhancing contextual reasoning is critical, as models struggle to weigh contextual factors when assessing dangers. The "Goldilocks effect"—where mid-sized models sometimes outperform larger ones—challenges the assumption that scaling parameters automatically improves safety performance. Category-specific optimization appears necessary; medical emergencies might benefit from specialized physiological feature detection, while disaster recognition could focus on environmental safety indicators. Contrastive learning with paired examples similar to our VERI dataset could help models better distinguish between visually similar but semantically distinct safety situations.
6 Related Work
For VLMs to function safely in real-world applications, accurate risk assessment is essential. While our research addresses everyday risk assessment across multiple domains, previous work has primarily focused on domain-specific applications.
In autonomous driving, Zhang et al. (2025) shows even larger models struggle with safety cognition, exhibiting limitations similar to our "overreaction problem." In healthcare, Ray et al. (2024)’s ERVQA dataset revealed medical visual reasoning challenges. Unlike ERVQA’s medical question-answering focus, our dataset addresses the fundamental challenge of distinguishing between genuine emergencies and visually similar safe scenarios across domains.
Fraser and Kiritchenko (2024)’s PAIRS dataset uses parallel images that differ only in demographic attributes to evaluate social biases in VLMs. While it focuses on bias detection, it shares our approach of using AI-generated contrastive image pairs with controlled variations. Similarly, Tu et al. (2023)’s benchmark evaluates VLMs’ tendency to hallucinate non-existent objects, representing another critical safety concern complementary to our overreaction problem.
Previous research has documented VLMs’ deficiencies in commonsense reasoning, with models failing at simple tasks like identifying a lemon from "tastes sour" (Ye et al., 2022), achieving only $<42\%$ performance compared to human performance of $83\%$. The digital twin modeling field offers a complementary perspective, where Yang et al. (2025) highlight how VLMs enable more flexible safety assessment through zero-shot learning. Recent evaluations show inconsistent scaling patterns (Zhang et al., 2025), aligning with our findings on the persistent "overreaction problem" across architectures and parameter scales.
7 Conclusion
We introduce VERI, a contrastive benchmark evaluating VLMs’ emergency recognition capabilities. Our findings reveal a systematic "overreaction problem" with false positive rates of 28-48% across models, primarily stemming from contextual overinterpretation (88-93% of errors). This bias persists regardless of model scale, with mid-sized models sometimes outperforming larger variants. These results highlight the need for targeted improvements in contextual reasoning rather than simply scaling model size, suggesting future research in specialized architectures and contrastive learning approaches for safety-critical visual assessment.
Limitations
Our binary classification approach (safe/dangerous) prevents exploration of severity gradations within emergency situations, which is an important aspect of real-world emergency assessment. In practice, emergency responses often depend on assessing degrees of risk rather than making binary determinations.
Despite our multi-stage creation process, the synthetic nature of our images may introduce subtle artifacts that differ from real-world photographs. These distinctions could potentially influence model performance in ways that might not generalize to natural images.
Our evaluation protocol simplifies the nuanced decision-making required in actual emergency situations, and our use of GPT-4o for judgment may not perfectly align with assessments from domain experts such as medical professionals or emergency responders.
Laboratory evaluations with controlled image pairs may not accurately predict performance in deployment scenarios with imperfect inputs or varying conditions. While our VERI dataset of 200 images (100 contrastive pairs) was deliberately designed as a focused diagnostic benchmark (similar to WSC with 60 examples), this approach necessarily makes trade-offs between depth and breadth, potentially limiting the generalizability across different cultural and geographical contexts.
Ethics Statement
We adhered to strict ethical guidelines in dataset creation, generating images with reasonable levels of danger without unnecessarily distressing content. Using AI-generated images rather than real emergency footage avoids ethical concerns related to exploitation of actual situations. Three independent annotators verified each image pair to ensure appropriate visual similarity and semantic distinction.
VERI is intended solely for research to improve safety-aware AI systems and should not replace human judgment in emergency response. We acknowledge that both false positives and negatives carry significant consequences—resource misallocation and delayed interventions, respectively. We advocate for balanced sensitivity and specificity in safety-critical systems with appropriate human oversight in any deployment.
Our findings about systematic biases in VLMs highlight the need for specialized training approaches rather than simply increasing model parameters to develop more contextually aware systems that correctly distinguish between genuine emergencies and visually similar but safe situations.
Acknowledgements
This research was supported by Brian Impact, a non-profit organization dedicated to advancing science and technology.
Appendix A Emergency Response Evaluation Details
This appendix provides additional details about our evaluation of emergency response quality (Q2), including the scoring protocol, representative model outputs, and category-specific analysis. These results complement our main findings by highlighting how models vary in their ability to generate accurate and context-sensitive responses once a danger is correctly identified.
A.1 GPT-4o Evaluation Methodology
To assess the quality of model-generated emergency responses (Q2), we used GPT-4o as an automatic evaluator. The prompt shown in Figure 7 was used to ensure consistency and transparency in scoring. This prompt provides the model with a description of the emergency situation, a gold-standard reference response, and the candidate response from the VLM. GPT-4o is then asked to assign a score between 0 and 1 based on specificity, factual accuracy, and alignment with expert protocols.
We additionally verified a random 10% subset of responses with two human annotators, achieving inter-rater agreement of $\kappa=0.83$, confirming that GPT-4o judgments were largely consistent with expert evaluations.
GPT-4o Evaluation Prompt The following is a description of an emergency situation in an image, a reference emergency response, and a response generated by a Vision-Language Model (VLM). - Image caption: "[CAPTION]" - Reference: "[REFERENCE_RESPONSE]" - VLM response: "[VLM_RESPONSE]" Evaluate how appropriate the VLM response is based on the situation and in comparison to the reference response. Your evaluation should strictly consider whether the VLM response aligns well in tone, intent, and direction with the reference response, is specific to the situation, and is factually accurate without misleading or overly vague content. Return only the following format: Score: <a number between 0 and 1> Reason: <brief explanation within 100 tokens>
Figure 7: The prompt used to evaluate emergency response quality with GPT-4o.
A.2 Category-Specific Performance Data
To better understand how emergency type affects model response quality, Table 5 presents Q2 scores disaggregated by category (PME, AB, ND) for all evaluated models.
The InternVL3 family exhibits increasing consistency across emergency categories as model size grows, with the 14B variant yielding nearly identical scores across all three categories (maximum deviation: 0.004). This suggests that larger InternVL3 models generalize better across diverse emergency types. In contrast, the Mistral family shows stronger category-specific preferences even at large scales, with Pixtral-Large performing markedly better on AB (0.70) and ND (0.70) than on PME (0.63), indicating potential limitations in medical response generalization.
Smaller models (<8B) exhibit larger performance differences across categories compared to larger models (>20B), with exceptions. Performance on Medical Emergencies shows the most consistent improvement with model size across families, suggesting that medical intervention knowledge particularly benefits from increased capacity. These patterns reinforce our main finding that emergency response (Q2) performance is more stable across categories than risk identification (Q1), but also highlight family-specific trends—such as Mistral models’ preference for AB and ND, and InternVL3’s balanced scaling.
Model PME AB ND
Qwen2.5-VL Family
Qwen2.5-VL (3B) 0.447 0.429 0.506
Qwen2.5-VL (7B) 0.633 0.589 0.606
Qwen2.5-VL (32B) 0.690 0.677 0.722
Qwen2.5-VL (72B) 0.661 0.744 0.681
LLaVA-Next Family
LLaVA-Next (7B) 0.466 0.500 0.433
LLaVA-Next (13B) 0.570 0.466 0.479
InternVL3 Family
InternVL3 (2B) 0.473 0.500 0.516
InternVL3 (8B) 0.569 0.615 0.626
InternVL3 (14B) 0.638 0.640 0.636
Mistral Family
Mistral-Small (24B) 0.622 0.609 0.644
Pixtral (12B) 0.545 0.633 0.588
Pixtral-Large (124B) 0.625 0.700 0.697
Other Models
Idefics2 (8B) 0.515 0.443 0.444
Phi-3.5-vision (4B) 0.500 0.462 0.471
Table 5: Emergency Response (Q2) scores across different emergency categories (PME: Personal Medical Emergencies, AB: Accidents & Behaviors, ND: Natural Disasters)
Emergency Response and Evaluation Examples Refer to caption [Accident & Behaviors Example: Burn Hazard]
Reference Response:
Immediately move the child away from the hot kettle… If the child has already touched the hot surface, run the affected area under cool water for 10-15 minutes. Do not apply ice, butter, or ointments… For serious burns with blistering, seek medical attention promptly…
High-Quality Model Response (Score: 0.9):
"…immediately remove the child from the hazardous situation. Gently but swiftly guide them away from the kettle… If burns are present, cool the affected area with lukewarm water and cover it with a sterile, non-adhesive bandage. Avoid ice or cold water, as it can worsen the injury…"
Low-Quality Model Response (Score: 0.2):
"…the little girl is playing with a kettle and an electric teapot. She is trying to turn on the kettle, but it is not working… The emergency response should be to check the power supply and make sure that the kettle and teapot are plugged in correctly…"
Refer to caption [Natural Disaster Example: Lightning Hazard]
Reference Response:
Surfers should immediately exit the water as lightning is extremely dangerous when combined with salt water… Follow the 30/30 rule: if thunder is heard within 30 seconds of seeing lightning, the storm is dangerously close. Exit the water quickly and move far from the shoreline toward substantial shelter… Wait at least 30 minutes after the last lightning strike before returning to the water…
High-Quality Model Response (Score: 0.9):
"…individuals surfing in what appears to be stormy weather, with lightning present… This situation poses immediate risks… **Immediate Evacuation**: Surfers should immediately exit the water and seek shelter as quickly as possible to avoid lightning strikes. **Avoid Tall Structures or Water**: Moving to a flat, low-lying area away from water, tall structures, or trees is crucial to minimize risk…" Low-Quality Model Response (Score: 0.2):
"…a group of people surfing in the ocean… they are facing a large wave that is about to crash. The emergency response should be to warn the people about the approaching wave and instruct them to paddle away from it… Additionally, it is essential to have a lifeguard or a rescue team nearby…"
Figure 8: Examples of high-quality versus low-quality emergency responses. High-quality responses provide specific, prioritized safety instructions aligned with expert protocols. Low-quality responses reveal critical failure modes, including misinterpretation of the actual threat (e.g., treating a burn hazard as an appliance malfunction, or focusing on wave height instead of lightning risk) and failure to recommend urgent, situation-specific actions.
A.3 Emergency Response and Evaluation
While correctly identifying emergencies is important, it is equally critical that models recommend appropriate and context-specific responses. Figure 8 presents representative cases from our evaluation. Each example includes a gold-standard reference response, a high-quality model output aligned with expert protocols, and a low-quality output that either misinterprets the situation or fails to provide actionable guidance. These contrasts reveal common patterns of strength and failure in VLMs’ emergency reasoning.
Contextual Overinterpretation Examples Refer to caption [Personal Medical Emergencies Example: Card Magician]
InternVL3 (8B):
"The situation appears dangerous because the bumper cars are colliding, which can lead to injuries from the impact or from the sparks flying, indicating a possible short circuit or mechanical failure."
LLaVa-Next (7B):
"The card could be sharp, and if the man were to accidentally bite down on it, he could cause himself injury. Additionally, there is a risk of the card being dislodged from his mouth and causing harm to others nearby. This situation is not safe and should be avoided."
Refer to caption [Accident & Behaviors Example: Farmer Harvesting Carrots]
Qwen2.5-VL (32B):
"The individuals in the image are not wearing any visible protective gear, such as gloves, safety goggles, or sturdy footwear. When working with soil, especially when using tools like the hand fork visible in the image, there is a risk of injury from sharp objects, splinters, or even hidden hazards like rocks or broken glass in the soil."
LLaVa-Next (7B):
"The situation appears dangerous because the man is in close proximity to the soil, which can be unstable and may cause him to lose his balance or be buried if the soil collapses. The environment also looks like it could be prone to sudden changes in weather or soil conditions, which could further increase the risk of accidents."
Figure 9: Examples of contextual overinterpretation in visually safe scenarios. The models infer safety risks that are either irrelevant or exaggerated given the visual evidence (e.g., treating a card trick as a choking hazard, or interpreting routine farming as a collapse risk). These cases illustrate how vision-language models may misattribute threats by overapplying safety heuristics in benign contexts.
Appendix B Detailed Contextual Overinterpretation Example
Our main analysis revealed that Contextual Overinterpretation accounts for 88-93% of model misclassifications across evaluated VLMs. This section provides additional examples of this systematic error pattern, illustrating how models exaggerate risks in safe scenarios across different categories. As shown in Figure 9, models frequently misinterpret harmless activities—from card tricks and eating spaghetti to gardening—as dangerous situations requiring intervention. These examples demonstrate how VLMs can correctly identify visual elements but consistently fail to assess their contextual safety implications, revealing a persistent "better-safe-than-sorry" bias that manifests across different visual domains and model architectures.
Category Total Errors CO % VM %
Accidents & Behaviors 326 85.3% 14.7%
Natural Disasters 298 100.0% 0.0%
Personal Medical 237 85.7% 14.3%
Table 6: Distribution of error types by emergency category. CO: Contextual Overinterpretation, VM: Visual Misinterpretation. Note that the Natural Disasters category exhibits exclusively Contextual Overinterpretation errors.
Model Total Errors CO % VM %
Qwen2.5-VL Family
Qwen2.5-VL (3B) 96 92.7% 7.3%
Qwen2.5-VL (7B) 71 91.5% 8.5%
Qwen2.5-VL (32B) 62 90.3% 9.7%
Qwen2.5-VL (72B) 48 91.7% 8.3%
LLaVA-Next Family
LLaVA-Next (7B) 71 88.7% 11.3%
LLaVA-Next (13B) 74 90.5% 9.5%
InternVL3 Family
InternVL3 (2B) 55 89.1% 10.9%
InternVL3 (8B) 31 90.3% 9.7%
InternVL3 (14B) 50 88.0% 12.0%
Mistral Family
Mistral-Small (24B) 71 91.5% 8.5%
Pixtral (12B) 47 89.4% 10.6%
Pixtral-Large (124B) 57 93.0% 7.0%
Other Models
Idefics2 (8B) 85 89.4% 10.6%
Phi-3.5-vision (4B) 43 88.4% 11.6%
Table 7: Distribution of error types across all evaluated models. CO: Contextual Overinterpretation, VM: Visual Misinterpretation.
Appendix C Detailed Error Pattern Analysis
Our in-depth analysis of risk identification (Q1) errors revealed a remarkably consistent distribution across both models and categories. Regardless of architecture or parameter count (2B–124B), all evaluated models exhibited a strong bias toward Contextual Overinterpretation (CO), which accounted for 88–93% of false positives.
This trend held across model sizes: CO rates were uniformly high across scale groups—91.4% for 0–5B models, 90.2% for 5–10B, 89.5% for 10–20B, and 90.9% for models above 20B. Such consistency suggests that limitations in contextual reasoning are systemic within current VLM architectures and cannot be resolved by scaling alone. Even the top-performing model in terms of precision (InternVL3-8B) misclassified 90.3% of its false positives due to contextual overinterpretation, while Pixtral-Large (124B)—the largest model—had an even higher CO rate of 93.0%. Table 7 summarizes CO/VM distributions across all models.
Category-level analysis (Table 6) further supports this pattern. Natural Disasters exhibited exclusively CO errors (100%), indicating that models recognize elements like fire, smoke, or water but fail to reason about containment or safety context. Similar but slightly more diverse error profiles were observed in Accidents & Behaviors (85.3% CO) and Personal Medical Emergencies (85.7% CO), where CO errors still dominated but Visual Misinterpretations (VM) occasionally occurred.
Taken together, these findings suggest that while VLMs can detect visual features associated with danger, they struggle to weigh contextual cues accurately—particularly in ambiguous or representational scenarios.